Calculate the voltage in the capacitor and the current in the inductor of the circuit, calculate the reactive power (Q), active (P) of the circuit and the total apparent power (S), which power factor is in the circuit.

0.5 H M 50 cos(10+ 30°) 6 1012 Z +v
Answer #1

0.5H = jWL: 15 V: 50 cos (10t+30) 102 • 50/30° Ž Zor I No 2 Ixyox tos -j2 50 L30 20 T VO . By : 10*5 = (2+40) 2 -j2 Vo: 50/36S: V¥ 1 (0) = 50 L30º y 17.677 /1050 35-355 136 * 12.5 Liosº 15=44193 LI35° VA) S: -312.49 +312:49 1 P= 312:49 W 0 = 312-49 V